# Thermodynamic and energetic constraints on transition probabilities of small-scale quantum systems

### v2

@@ I Introduction
Indeed, in particular for nanoelectronics systems, fluctuation-dissipation bounds have recently been developed [43] for current fluctuations (or noise) in the presence of a temperature bias. Although these bounds apply to conductors with generic transmission properties, their validity is limited to systems with weak electron-electron interactions. In the presence of possibly strong interactions, the perturbative approach developed in Refs. [25, 26] showed that noise is super-Poissonian in the tunneling regime, even in the presence of a temperature bias. However, the constraint is rather loose and the role of the temperature bias is not singled out. Hence, there remains the important question of whether fundamental bounds on the dynamics for strongly interacting systems exist accounting for the impact of a temperature bias.
@@ II.1 Transition probabilities
$S(\beta)$ $\equiv-\text{Tr}\left\{\frac{e^{-\beta\hat{H}_{0}}}{Z(\beta)}\log\frac{e^{-\beta\hat{H}_{0}}}{Z(\beta)}\right\}.$ (5b)
@@ II.2 Thermodynamic constraint
$\tilde{\epsilon}=\frac{1}{\beta_{\text{c}}-\beta_{\text{h}}}\log\frac{Z(\beta_{\text{h}})}{Z(\beta_{\text{c}})}.$ (6)
@@
for all $\epsilon_{a}$ and $\beta_{\text{c}}>\beta_{\text{h}}$. Summing (7) over all eigenstates $a$ leads to the statement that the internal energy is an increasing function of temperature, $U(\beta_{\text{h}})\geq U(\beta_{\text{c}})$. We furthermore note that the contributions of (7) can be rewritten in terms of a response of the probabilities to a temperature variation
@@
This thermodynamic constraint (12) on the transition probabilities at different temperatures is the first main result of this paper. It implies that the response of the transition probabilities to a change in temperature is limited by the transition probabilities themselves and by the thermodynamic cost of changing the system’s temperature. In order to intuitively understand how far it constrains the system dynamics at different initial temperatures, we consider the thermodynamics of cooling down a hot system in order to be able to subsequently perform the two-point measurement scheme starting from a cold state, as sketched in Fig. 2. Analogously, we consider the thermodynamics of heating up a cold system in order to be able to subsequently perform the two-point measurement scheme starting from a hot state, as sketched in Fig. 2.
@@
    Heating: The system initially at $T_{\text{c}}$ is heated up by bringing it into contact with a bath at $T_{\text{h}}<T_{\text{c}}$ [Fig. 2(b)]. Heat flows into the system until it reaches temperature $T_{\text{h}}$, thus inducing the change $[U(\beta_{\text{h}})-U(\beta_{\text{c}})]$ in internal energy and $[S(\beta_{\text{h}})-S(\beta_{\text{c}})]$ in entropy. While heat is flowing, it is possible to extract work, which is at most $-\Delta F^{\text{h}}$. After the system has been heated up, the two-point measurement scheme is performed and leads to the transition probabilities $p_{a\to b}(\beta_{\text{h}})$.
Furthermore, the factors $\frac{\beta_{\text{c}}}{\beta_{\text{c}}-\beta_{\text{h}}}=\frac{T_{\text{h}}}{T_{\text{h}}-T_{\text{c}}}$ and $\frac{\beta_{\text{h}}}{\beta_{\text{c}}-\beta_{\text{h}}}=\frac{T_{\text{c}}}{T_{\text{h}}-T_{\text{c}}}$ in Eq. (14) correspond to the coefficient of performance of a heat pump and of a refrigerator, respectively. Consequently, the product $-\frac{\beta_{\text{c}}\Delta F^{\text{c}}}{\beta_{\text{c}}-\beta_{\text{h}}}$ sets a lower limit on the heat absorbed by the cold bath during the cooling process when the extracted work is maximum, as detailed in Appendix A. Similarly, the product $-\frac{\beta_{\text{h}}\Delta F^{\text{h}}}{\beta_{\text{c}}-\beta_{\text{h}}}$ sets a lower limit on the energy absorbed by the system during the heating process. Thus, in the thermodynamic cost function of Eq. (14) the hot and cold transition probabilities are weighted by the heat dissipated to cool down or heat up the system, respectively.
@@ II.3 Energetic constraint
II.4 Thermodynamic limit
Here, we comment on the relevance of the constraints (12) and (17) when the system approaches the thermodynamic limit. Notably, both the thermodynamic and the energetic constraints combine transition probabilities which are always smaller than one, with extensive properties of the system, namely with its internal energy and with its nonequilibrium free energy. However, the extensive quantities only appear on the left-hand side of Eqs. (12, 17), and not on the right-hand side. This feature implies that the left- and the right-hand sides behave very differently depending on the size of the system. More concretely, if the extensive quantities in the thermodynamic limit scale as
where $\lambda$ is the scale parameter, and the transition probabilities scale with an arbitrary scaling function $f(\lambda)$, i.e. $p_{a\to b}\to f(\lambda)p_{a\to b}$, both the thermodynamic constraint of Eq. (12) and the energetic constraint of Eq. (17) become trivial. Indeed, for $\lambda\to\infty$, the right-hand side containing the response $\mathcal{W}^{(\text{Resp})}_{a\to b}$ becomes negligible, and Eqs. (12, 17) then reduce to expressions that state the positivity of the thermodynamic and energetic costs, respectively, on the left-hand side.
However, the fact that the right hand sides of these equations can be neglected in the thermodynamic limit—independently on how it is taken—leads to a trivial statement. The thermodynamic and energetic costs are positive by construction: The transition probabilities are positive $p_{a\to b}\geq 0$, and so are the coefficients of performance $\eta^{\text{(c,h)}}\geq 0$, the nonequilibrium free energies $-\Delta F^{\text{(c,h)}}\geq 0$, as well as the energy variation $dU(x)$ in the integral in Eq. (17). Thus, both the thermodynamic and the energetic constraints pose constraints on small-scale quantum systems that do not satisfy the thermodynamic limit.
II.5 Saturating constraints and trivial constraints
@@
In contrast, the constraints are always trivially fulfilled when the function characterizing the response to a temperature variation, $\mathcal{W}^{\text{(Resp)}}_{a\to b}$ is negative. This is the case, when the energy of the initial state $\epsilon_{a}$ differs from the internal energy of the system by more than the internal energy’s standard deviation, see Appendix B. This means that the constraints are nontrivial only for those transition rates, where the initial state is “typical” for the system, namely with an energy close to the internal energy of the system.
@@ III.1 Weak tunnel coupling
which we assume to be a small perturbation, and which induces transitions in the subsystems. Further, in the weak tunnel-coupling regime, we describe the total system with the product state $\hat{\rho}=\hat{\rho}^{\mathrm{L}}\otimes\hat{\rho}^{\mathrm{R}}$, where $\hat{\rho}^{\alpha}$ is the Gibbs state, defined in Eq. (3), at inverse temperature $\beta_{\alpha}$. With this, the rates $\Gamma_{\rightleftarrows}$ for absorbing or emitting a quantum of energy $\hbar\omega$, induced by the coupling, are—in the spectral representation [18]—given as
@@
Interestingly, the transition rates in the weak tunneling regime can be directly connected to transport quantities, namely to a current and its zero-frequency noise [18, 25]
@@ III.2 Constraints in the tunneling regime
$\mathcal{W}^{\text{(Thermo)}}_{\rightleftarrows}$$\equiv$ $-\Delta F_{\text{L}}^{\text{(c)}}\eta^{\text{(h)}}\Gamma_{\rightleftarrows}(\beta_{\mathrm{h}},\beta_{\mathrm{c}})+$ (24b)
@@
We emphasize that Eq. (23a) directly translates into a constraint on the out-of-equilibrium noise, which is given by the sum of the tunneling rates via Eq. (22b). This also means, from a more practical, experimental viewpoint, that one possibility to test the constraint (23a) is by measuring the noise and tunneling current in two configurations: (i) When both subsystems have the same cold temperature, and (ii) in the desired out-of-equilibrium condition. One can then access the tunneling rates as $2q^{2}\Gamma_{\rightleftarrows}=\mathcal{S}\pm qI$. All the other quantities entering the inequality can be calculated once the state of the subsystem considered at the different temperatures is known  [^2] Note that it is not necessary to know the Hamiltonian of the subsystem kept at fixed temperature, which may be arbitrarily complicated, to test these constraints.. Similarly to the thermodynamic constraint, Eq. (23b) directly translates into an energetic constraint on the integral over the out-of-equilibrium noise in Eq. (22b) by taking the sum of the energetic constraints for both tunneling directions. However, this integral makes the energetic constraint less relevant from an experimental point of view because it requires knowledge of both internal energy and tunneling rates at all intermediate temperatures of subsystem L.
From the thermodynamic constraint (23a), we can also derive a direct lower bound on the out-of-equilibrium tunneling rates $\Gamma_{\rightleftarrows}(\beta_{\mathrm{h}},\beta_{\mathrm{c}})$ in terms of the equilibrium rates (see the derivation in Appendix E). These constraints furthermore contain the rate response, (integrals of) thermodynamic quantities, and they read
@@ III.3 Atom coupled to nonlinear cavity
Here, one of the subsystems is the atom, described by the Pauli matrix $\hat{\sigma}_{z}$ and characterized by the frequency $\omega_{\text{A}}$. The cavity is the other subsystem, and is described by the number operator $\hat{n}=\hat{a}^{\dagger}\hat{a}$, with $[\hat{a},\hat{a}^{\dagger}]=1$, and the cavity frequency $\omega_{\text{C}}$. We also include a Kerr nonlinearity, parametrized by $u_{\text{C}}$, which plays the role of effective interactions between cavity photons in this system. The atom and the cavity exchange photon quanta through the weak tunneling term $g(\hat{a}\hat{\sigma}_{+}+\hat{a}^{\dagger}\hat{\sigma}_{-})$ with $g\ll\hbar\omega_{\text{A,C}}$ and where $\hat{\sigma}_{+}$ and $\hat{\sigma}_{-}$ are the raising and lowering operators of the atom states. Note that, in the language of Eq. (20), we choose $\omega=0$ since any external driving frequency can be incorporated in the laser detuning—replacing $\omega_{\mathrm{C}}$—in the rotating frame [41]. For $u_{\text{C}}=0$, the Hamiltonian (30) reduces to the Jaynes-Cummings Hamiltonian [58].
@@ III.4 Coupled fermionic rings
Figure 6: Thermodynamic cost (dashed lines), energetic cost (dotted lines), and rate response (solid lines) for two (U and D) coupled fermionic rings with sizes $L_{\text{D}}=3$ and $L_{\text{U}}=10$. The costs are plotted vs the inverse temperature difference $\beta_{\text{D}}-\beta_{\text{U}}$. The ring hopping parameters, $t_{\text{D}}$ and $t_{\text{U}}$, and the driving frequency are fixed as $t_{\text{U}}=\frac{4}{5}t_{\text{D}}=2\hbar\omega$. The upper ring charging energy $u_{\text{U}}=0$ is also fixed, while the down ring charging energy is taken as $u_{\text{D}}/2t_{\text{D}}=0,10$ for curves in blue and red, respectively. In panels (a, c) the upper, larger ring is kept at the cold temperature ($\beta_{\text{U}}>\beta_{\text{D}}$), whereas in panels (b, d) it instead is the down, smaller ring that is colder ($\beta_{\text{L}}>\beta_{\text{U}}$).
@@
Similarly to the atom-cavity system considered in Sec. III.3, the charging energy $u_{\text{D}}$ influences the thermodynamic and energetic constraints in the system of coupled rings. In panels (a, c) of Fig. 6, we show results where the lower ring, namely the one displaying finite interaction, is the one that is considered at two different temperatures when comparing out-of-equilibrium rates. This produces nontrivial constraints for both $\Gamma_{\rightleftarrows}$, similar to the case of the atom-cavity system, where the cavity with nonlinearity $u_{\mathrm{C}}$ is taken at two different temperatures, as shown in Fig. 5 (b, d). In contrast, as shown in panels (b, d) of Fig. 6, for the case where the upper ring is taken at two different temperatures only, the rate $\Gamma_{\leftarrow}$ is non-trivially constrained, as long as the charging energy $u_{\text{U}}=0$ vanishes.
Furthermore, as discussed in Sec. II.4, increasing the size of the ring that is considered at two different temperatures (namely corresponding to the one on which the two-point measurement is performed, see Sec. II), when comparing out-of-equilibrium rates, weakens the constraints, as is seen in panels (b, d) of Fig. 7. In this example, the tunneling rates scale inversely with the ring size, i.e. with a scaling function $f$ on the form $f(L_{\text{U}})\sim 1/L_{\text{U}}$ [see below Eq. (19)], whereas the extensive quantities entering the thermodynamic and energetic costs scale linearly with $L_{\text{U}}$. Thus, while the costs are essentially unaffected upon increasing $L_{\text{U}}$, the rate responses instead decrease. By contrast, if the ring that is considered at two different temperatures has a fixed size, both costs and rate responses scale as $1/L_{\text{U}}$ because of the scaling of the tunneling rates alone, see panels (a, c). This feature implies that, as long as the size of the subsystem considered at two different temperatures remains small, there are nontrivial constraints on the tunneling rates, irrespective of the size of the subsystem with fixed temperature.

### v3

@@ I Introduction
Indeed, in particular for nanoelectronic systems, fluctuation-dissipation bounds have recently been developed [43] for current fluctuations (or noise) in the presence of a temperature bias. Although these bounds apply to conductors with generic transmission properties, their validity is limited to systems with weak electron-electron interactions. In the presence of possibly strong interactions, the perturbative approach developed in Refs. [25, 26] showed that noise is super-Poissonian in the tunneling regime, even in the presence of a temperature bias. However, this constraint does not single out the role of the temperature bias. Hence, there remains the important question of whether fundamental bounds on the dynamics for strongly interacting systems exist accounting for the impact of a temperature bias.
@@ II.1 Transition probabilities
$S(\beta)$ $\equiv-\text{Tr}\left\{\frac{e^{-\beta\hat{H}_{0}}}{Z(\beta)}\ln\frac{e^{-\beta\hat{H}_{0}}}{Z(\beta)}\right\}.$ (5b)
@@ II.2 Thermodynamic constraint
$\tilde{\epsilon}=\frac{1}{\beta_{\text{c}}-\beta_{\text{h}}}\ln\frac{Z(\beta_{\text{h}})}{Z(\beta_{\text{c}})}.$ (6)
@@
for all $\epsilon_{a}$ and $\beta_{\text{c}}>\beta_{\text{h}}$. Summing (7) over all eigenstates $a$ leads to the statement that the internal energy is an increasing function of temperature, $U(\beta_{\text{h}})\geq U(\beta_{\text{c}})$. We furthermore note that the contributions of (7) can partially be rewritten in terms of a response of the probabilities to a temperature variation
@@
This thermodynamic constraint (12) on the transition probabilities at different temperatures is the first main result of this paper. It implies that the response of the transition probabilities to a change in temperature is limited by the transition probabilities themselves and by the thermodynamic cost of changing the system’s temperature. In order to intuitively understand how far it constrains the system dynamics at different initial temperatures, we consider the thermodynamics of cooling down a hot system in order to be able to subsequently perform the two-point measurement scheme starting from a cold state, as sketched in Fig. 2(a). Analogously, we consider the thermodynamics of heating up a cold system in order to be able to subsequently perform the two-point measurement scheme starting from a hot state, as sketched in Fig. 2(b).
@@
    Heating: The system initially at $T_{\text{c}}$ is heated up by bringing it into contact with a bath at $T_{\text{h}}>T_{\text{c}}$ [Fig. 2(b)]. Heat flows into the system until it reaches temperature $T_{\text{h}}$, thus inducing the change $[U(\beta_{\text{h}})-U(\beta_{\text{c}})]$ in internal energy and $[S(\beta_{\text{h}})-S(\beta_{\text{c}})]$ in entropy. While heat is flowing, it is possible to extract work, which is at most $-\Delta F^{\text{h}}$. After the system has been heated up, the two-point measurement scheme is performed and leads to the transition probabilities $p_{a\to b}(\beta_{\text{h}})$.
Furthermore, the factors $\eta^{(\mathrm{h})}\equiv\frac{\beta_{\text{c}}}{\beta_{\text{c}}-\beta_{\text{h}}}=\frac{T_{\text{h}}}{T_{\text{h}}-T_{\text{c}}}$ and $\eta^{(\mathrm{c})}\equiv\frac{\beta_{\text{h}}}{\beta_{\text{c}}-\beta_{\text{h}}}=\frac{T_{\text{c}}}{T_{\text{h}}-T_{\text{c}}}$ in Eq. (14) correspond to the coefficient of performance of a heat pump and of a refrigerator, respectively. Consequently, the product $-\frac{\beta_{\text{c}}\Delta F^{\text{c}}}{\beta_{\text{c}}-\beta_{\text{h}}}$ sets a lower limit on the heat absorbed by the cold bath during the cooling process when the extracted work is maximum, as detailed in Appendix A. Similarly, the product $-\frac{\beta_{\text{h}}\Delta F^{\text{h}}}{\beta_{\text{c}}-\beta_{\text{h}}}$ sets a lower limit on the energy absorbed by the system during the heating process. Thus, in the thermodynamic cost function of Eq. (14) the hot and cold transition probabilities are weighted by the heat dissipated to cool down or heat up the system, respectively.
@@ II.3 Energetic constraint
II.4 Saturating constraints and trivial constraints
@@
In contrast, the constraints are always trivially fulfilled when the function characterizing the response to a temperature variation, $\mathcal{W}^{\text{(Resp)}}_{a\to b}$ is negative, because the theremodynamic and energetic costs are positive by construction. This is the case, when the energy of the initial state $\epsilon_{a}$ differs from the internal energy of the system by more than the internal energy’s standard deviation, see Appendix B. This means that the constraints are nontrivial only for those transition rates, where the initial state is “typical” for the system, namely with an energy close to the internal energy of the system.
II.5 Thermodynamic limit
Here, we comment on the relevance of the constraints (12) and (17) when the system approaches the thermodynamic limit. Notably, both the thermodynamic and the energetic constraints combine transition probabilities which are always smaller than one, with extensive properties of the system, namely with its internal energy and with its nonequilibrium free energy. However, the extensive quantities only appear explicitly on the left-hand side of Eqs. (12, 17). This feature implies that the left- and the right-hand sides behave very differently depending on the size of the system. More concretely, if the extensive quantities in the thermodynamic limit scale as
where $\lambda$ is the scale parameter, and the transition probabilities scale with an arbitrary scaling function $f(\lambda)$, i.e. $p_{a\to b}\to f(\lambda)p_{a\to b}$, both the thermodynamic constraint of Eq. (12) and the energetic constraint of Eq. (17) become trivial. Indeed, the right-hand side contribution from the response $\mathcal{W}^{(\text{Resp})}_{a\to b}$ is smaller than the standard deviation of the internal energy, as shown in Appendix B. Therefore, for $\lambda\to\infty$, the right-hand side becomes negligible and Eqs. (12, 17) then reduce to expressions that state the positivity of the thermodynamic and energetic costs, respectively, on the left-hand side.
However, the fact that the right-hand sides of these equations can be neglected in the thermodynamic limit—independently on how it is taken—leads to a trivial statement. The thermodynamic and energetic costs are positive by construction: The transition probabilities are positive $p_{a\to b}\geq 0$, and so are the coefficients of performance $\eta^{\text{(c,h)}}\geq 0$, the nonequilibrium free energies $-\Delta F^{\text{(c,h)}}\geq 0$, as well as the energy variation $dU(x)$ in the integral in Eq. (17). Thus, both the thermodynamic and the energetic constraints pose constraints on small-scale quantum systems that do not satisfy the thermodynamic limit.
@@ III.1 Weak tunnel coupling
which we assume to be a small perturbation, and which induces transitions in the subsystems. Further, in the weak tunnel-coupling regime, we describe the total system with the product state $\hat{\rho}=\hat{\rho}^{\mathrm{L}}\otimes\hat{\rho}^{\mathrm{R}}$, where $\hat{\rho}^{\alpha}$ is the Gibbs state, defined in Eq. (3), at inverse temperature $\beta_{\alpha}$. With this, the rates $\Gamma_{\rightleftarrows}$ for absorbing or emitting a quantum of energy $\hbar\omega$, induced by the coupling, are—in the spectral representation [18, 24]—given as
@@
Interestingly, the transition rates in the weak-tunneling regime can be directly connected to transport quantities, namely to a current and its zero-frequency noise [18, 25]
@@ III.2 Constraints in the tunneling regime
$\mathcal{W}^{\text{(Thermo)}}_{\rightleftarrows}$$\equiv$ $-\Delta F_{\text{L}}^{\text{(c)}}\eta^{\text{(h)}}\Gamma_{\rightleftarrows}(\beta_{\mathrm{h}},\beta_{\mathrm{c}})$ (24b)
@@
We emphasize that Eq. (23a) directly translates into a constraint on the out-of-equilibrium noise, which is given by the sum of the tunneling rates via Eq. (22b). This also means, from a more practical, experimental viewpoint, that one possibility to test the constraint (23a) is by measuring the noise and tunneling current in two configurations: (i) When both subsystems have the same cold temperature, and (ii) in the desired out-of-equilibrium condition. One can then access the tunneling rates as $2q^{2}\Gamma_{\rightleftarrows}=\mathcal{S}\pm qI$. All the other quantities entering the inequality can be calculated once the state of the subsystem considered at the different temperatures is known  [^2] Note that it is not necessary to know the Hamiltonian of the subsystem R kept at fixed temperature, which may be arbitrarily complicated, to test these constraints.. Similarly to the thermodynamic constraint, Eq. (23b) directly translates into an energetic constraint on the integral over the out-of-equilibrium noise in Eq. (22b) by taking the sum of the energetic constraints for both tunneling directions. However, this integral makes the energetic constraint less relevant from an experimental point of view because it requires knowledge of both internal energy and tunneling rates at all intermediate temperatures of subsystem L.
From the thermodynamic constraint (23a), we can also derive a direct lower bound on the out-of-equilibrium tunneling rates $\Gamma_{\rightleftarrows}(\beta_{\mathrm{h}},\beta_{\mathrm{c}})$ in terms of the equilibrium rates (see the derivation in Appendix E). These constraints furthermore contain the rate response, (integrals) of thermodynamic quantities, and they read
@@ III.3 Atom coupled to nonlinear cavity
Here, one of the subsystems is the atom, described by the Pauli matrix $\hat{\sigma}_{z}$ and characterized by the frequency $\omega_{\text{A}}$. The cavity is the other subsystem, and is described by the number operator $\hat{n}=\hat{a}^{\dagger}\hat{a}$, with $[\hat{a},\hat{a}^{\dagger}]=1$, and the cavity frequency $\omega_{\text{C}}$. We also include a Kerr nonlinearity, parametrized by $u_{\text{C}}$, which plays the role of effective interactions between cavity photons in this system. The atom and the cavity exchange photon quanta through the weak-tunneling term $g(\hat{a}\hat{\sigma}_{+}+\hat{a}^{\dagger}\hat{\sigma}_{-})$ with $g\ll\hbar\omega_{\text{A,C}}$ and where $\hat{\sigma}_{+}$ and $\hat{\sigma}_{-}$ are the raising and lowering operators of the atom states. Note that, in the language of Eq. (20), we choose $\omega=0$ since any external driving frequency can be incorporated in the laser detuning—replacing $\omega_{\mathrm{C}}$—in the rotating frame [41]. For $u_{\text{C}}=0$, the Hamiltonian (30) reduces to the Jaynes-Cummings Hamiltonian [58].
@@ III.4 Coupled fermionic rings
Figure 6: Thermodynamic cost (dashed lines), energetic cost (dotted lines), and rate response (solid lines) for two (U and D) coupled fermionic rings with sizes $L_{\text{D}}=3$ and $L_{\text{U}}=10$. The costs are plotted vs the inverse temperature difference $\beta_{\text{D}}-\beta_{\text{U}}$. The ring hopping parameters, $t_{\text{D}}$ and $t_{\text{U}}$, and the driving frequency are fixed as $t_{\text{U}}=\frac{4}{5}t_{\text{D}}=2\hbar\omega$. The upper ring charging energy $u_{\text{U}}=0$ is also fixed, while the down ring charging energy is taken as $u_{\text{D}}/2t_{\text{D}}=0,10$ for curves in blue and red, respectively. In panels (a, c) the upper, larger ring is kept at the cold temperature ($\beta_{\text{U}}>\beta_{\text{D}}$), whereas in panels (b, d) it instead is the down, smaller ring that is colder ($\beta_{\text{D}}>\beta_{\text{U}}$).
@@
Similarly to the atom-cavity system considered in Sec. III.3, the charging energy $u_{\text{D}}$ influences the thermodynamic and energetic constraints in the system of coupled rings. In panels (a, c) of Fig. 6, we show results where the lower ring, namely the one displaying finite interaction, is the one that is considered at two different temperatures when comparing out-of-equilibrium rates. This produces nontrivial constraints for both $\Gamma_{\rightleftarrows}$, similar to the case of the atom-cavity system, where the cavity with nonlinearity $u_{\mathrm{C}}$ is taken at two different temperatures, as shown in Fig. 5 (b, d). In contrast, as shown in panels (b, d) of Fig. 6, for the case where the upper ring is taken at two different temperatures, only the rate $\Gamma_{\leftarrow}$ is non-trivially constrained, as long as the charging energy $u_{\text{U}}=0$ vanishes.
Furthermore, as discussed in Sec. II.5, increasing the size of the ring that is considered at two different temperatures (namely corresponding to the one on which the two-point measurement is performed, see Sec. II), when comparing out-of-equilibrium rates, weakens the constraints, as is seen in panels (b, d) of Fig. 7. In this example, the tunneling rates scale inversely with the ring size, i.e. with a scaling function $f$ on the form $f(L_{\text{U}})\sim 1/L_{\text{U}}$ [see below Eq. (19)], whereas the extensive quantities entering the thermodynamic and energetic costs scale linearly with $L_{\text{U}}$. Thus, while the costs are essentially unaffected upon increasing $L_{\text{U}}$, the rate responses instead decrease. By contrast, if the ring that is considered at two different temperatures has a fixed size, both costs and rate responses scale as $1/L_{\text{U}}$ because of the scaling of the tunneling rates alone, see panels (a, c). This feature implies that, as long as the size of the subsystem considered at two different temperatures remains small, there are nontrivial constraints on the tunneling rates, irrespective of the size of the subsystem with fixed temperature.